OpenStudy (anonymous):

Solving for equations by substitution . 15) 6x+6y=-6 5x +y = -13 My answers Please check! 6x+6y=-6 y=-5x-13 6x + 6(-5x -13) = -6 6x -30x -78 =-6 -24x-78+78 = -6+78 -24x=-72 -2x/-24= -72/-24 x= 3.25 6(3.25)+6y=-6 19.5+6y = -6 -19.5 -19.5 6y=-25.5 6y/6=-25.5/6 y = -4.25

OpenStudy (donblue22):

That seems like the right answers

OpenStudy (anonymous):

That 2 should be a 24 lol

OpenStudy (donblue22):

And just so you know a good way to check is plugging the answers into one of the original equations and seeing if it works

OpenStudy (anonymous):

I know lol. just need to check with others.

OpenStudy (anonymous):

no right answer is x = -3 and y=2

OpenStudy (anonymous):

Thank you! Can you show me where I went wrong.?

OpenStudy (anonymous):

-24x-78+78 = -6+78 it should be in this step -24x - 78 = -6

OpenStudy (anonymous):

why did you repeat 78 again it totally changed the equation

OpenStudy (nincompoop):

to isolate -24x and if you added the same amount on the right-hand side the equation remains the same

OpenStudy (anonymous):

I put it like that to show that I was moving it over to the other side of the equal sign

OpenStudy (anonymous):

but you dont need to remove it you just have a remaining last step which is -24x=-6 +78 .... x=-3

OpenStudy (nincompoop):

your division is the incorrect part you can check if you did it correctly by using the values of x and y into any of the two original equations